OpenStudy (skullpatrol):

Solve $$ax-a^2=bx-b^2$$ for $$x$$.

OpenStudy (ivettef365):

you need to isolate the x so the first thing you do is combine the x's as follows: ax - bx = a^2 -b^2

OpenStudy (ivettef365):

now take the x out and you have x(a-b) = a^2 =- b^2

OpenStudy (ivettef365):

then divide by a-b and you have x = a^2-b^2 --------- a - b

OpenStudy (anonymous):

x= (a-b) (a+b) divded by a-b X= a+b

OpenStudy (ivettef365):

right you can simplify

OpenStudy (anonymous):

x= a+b ..is the answer

OpenStudy (ivettef365):

correct

OpenStudy (skullpatrol):

Is that the "complete" answer?

OpenStudy (anonymous):

yes u cant simplify further

OpenStudy (skullpatrol):

Are you sure that is the ENTIRE answer???

OpenStudy (ivettef365):

what else do you want on the answer

OpenStudy (skullpatrol):

My book gives a restriction on x=a+b.

OpenStudy (ivettef365):

then don't simplify it, use the one I gave you before simplifying.

OpenStudy (skullpatrol):

But I will lose marks for not simplifying?

OpenStudy (ivettef365):

then that makes no sense, if you simplify that is the answer x= a+b, there is nothing else to it.

OpenStudy (skullpatrol):

$$a \neq b$$

OpenStudy (loser66):

\[a \neq b\] is condition?

OpenStudy (skullpatrol):

Yes, a restriction on the simplified answer x=a+b.

OpenStudy (loser66):

nope, my question is a not equal b is a restriction, too? I solved and got a=b then for all x you have that form

OpenStudy (loser66):

I mean you have infinite solution except x =a +b

OpenStudy (skullpatrol):

x=a+b is the solution; except when a=b, then there is no solution.

OpenStudy (loser66):

how can it be. ok, correct me if I am wrong, please

OpenStudy (loser66):

\[ax-a^2 =bx-b^2\\ax-a^2-bx+b^2=0 \] only one way to move terms without making any condition.

OpenStudy (loser66):

factor out to get (a-b) (x-(a+b)) =0

OpenStudy (loser66):

if you restrict the second part, so, the product =0 if and only if a=b

OpenStudy (loser66):

So , when a= b replace to original problem \[ax -a^2 = ax -a^2\] for all x

OpenStudy (loser66):

Point out my mistake, please

OpenStudy (skullpatrol):

$$ (a-b) (x-(a+b)) =0$$becomes a true statement when $$a-b=0$$or $$x-(a+b)=0$$ If $$a-b=0,$$then the solution set is {real numbers} If $$x-(a+b)=0,$$ then the solution set is all the values of x for which x =a+b